# Policy Contrastive Decoding for Robotic Foundation Models

### v1

@@ Abstract
Robotic foundation models, or generalist robot policies, hold immense potential to enable flexible, general-purpose and dexterous robotic systems. Despite their advancements, our empirical experiments reveal that existing robot policies are prone to learning spurious correlations from pre-training trajectories, adversely affecting their generalization capabilities beyond the training data. To tackle this, we propose a novel Policy Contrastive Decoding (PCD) approach, which redirects the robot policy’s focus toward object-relevant visual clues by contrasting action probability distributions derived from original and object-masked visual inputs. As a training-free method, our PCD can be used as a plugin to improve different types of robot policies without needing to finetune or access model weights. We conduct extensive experiments on top of three open-source robot policies, including the autoregressive policy OpenVLA and the diffusion-based policies Octo and $\pi_{0}$. The obtained results in both simulation and real-world environments prove PCD’s flexibility and effectiveness, e.g., PCD enhances the state-of-the-art policy $\pi_{0}$ by 8% in the simulation environment and by 108% in the real-world environment. Code and demos are publicly available at: https://Koorye.github.io/proj/PCD.
@@ 1 Introduction
Figure 1: Robot policies tend to spuriously correlate task-irrelevant features with actions, compromising their ability to generalize to unseen scenarios. As observed, changing the light position from (a) to (b) and the drawer handle position from (a) to (c) results in 36% and 32% drops in the performance of the baseline policy OpenVLA [20], respectively. More results are in Appendix A.1.
@@
Contributions. Our contributions in this work are threefold. 1) We propose PCD, a simple, training-free, and easy-to-implement scheme to tackle the spurious correlation issue of robot policies. 2) PCD can be used as a plugin to enhance both autoregressive and diffusion-based policies, without requiring fine-tuning or access to pre-trained model weights. 3) Extensive experiments demonstrate the flexibility and effectiveness of PCD, consistently improving performance over three state-of-the-art robot policies across 15 manipulation tasks on multiple robotic platforms. We believe that our PCD provides a novel perspective for achieving robust adaptation of robot policies to downstream tasks.
@@ 3.2 Policy Contrastive Decoding (PCD)
where $C$ is a normalization constant and a larger value of $\alpha\geq 0$ indicates a stronger amplification of differences between the two distributions. Particularly, $\alpha=0$ recovers the baseline policy. In such a manner, the obtained $p_{t}^{*}$ amplifies model predictions on object-relevant features in $\boldsymbol{u}$, thus exhibiting insensitivity to spurious features in $\boldsymbol{v}$, as illustrated in Fig. 2.
@@ 4 Experiments
In this section, we seek to answer the following questions: {mdframed} [skipabove=4pt, innertopmargin=4pt, innerbottommargin=4pt] 1) Can PCD improve robot policies in both simulation and real-world environments?
@@ 4.4 Robustness to Various Kinds of Spurious Correlations
The central concept of our proposed PCD approach in this work is to overcome the adverse effects of spurious features on the decision-making of robot policies. It is crucial to investigate the following question: What types of spurious correlations can our PCD effectively address? To this end, we evaluate the performance of the OpenVLA [20] and $\pi_{0}$ [3] policies across a range of scenario variations in both simulation and real-world environments, including changes in spatial relationships, brightness, distractors, and table textures in visual observations, as illustrated in Fig. 5. The results shown in the figure lead to the following observations. 1) Consistent with the previous experimental results, our PCD significantly improves the performance of the baseline policies on both simulation and real-world tasks. 2) Both baseline policies experience substantial drops in performance when testing scenarios are modified. For example, altering the brightness of visual observations results in 48% and 75% decreases in success rates for the two policies, respectively. 3) Our proposed PCD approach consistently alleviates the performance degradation of the baseline policies across all unseen testing scenarios. Remarkably, after integrating PCD, the two baseline policies even performs better in 4/10 of unseen scenarios than in the original training scenarios. In summary, the results in the figure reveal that robot policies often form spurious correlations between task-irrelevant features and actions, limiting their generalization capacity beyond the training data distribution. Our proposed PCD approach can serve as a plugin to mitigate the adverse effects of various types of spurious correlations on the generalization of robot policies.
Figure 5: Performance of baseline policies integraged w/ or w/o our proposed PCD approach in unseen testing scenarios. OpenVLA [20] and $\pi_{0}$ [3] are used as baselines in simulation and real-world environments, respectively. PCD effectively mitigates the side effects of various types of spurious correlations, boosting the robot policy’s generalization to novel scenarios.

### v2

@@ Abstract
Robotic foundation models, or generalist robot policies, hold immense potential to enable flexible, general-purpose and dexterous robotic systems. Despite their advancements, our empirical experiments reveal that existing robot policies are prone to learning spurious correlations from pre-training trajectories, adversely affecting their generalization capabilities beyond the training data. To tackle this, we propose a novel Policy Contrastive Decoding (PCD) approach, which redirects the robot policy’s focus toward object-relevant visual clues by contrasting action probability distributions derived from original and object-masked visual inputs. As a training-free method, our PCD can be used as a plugin to improve different types of robot policies without needing to finetune or access model weights. We conduct extensive experiments on top of three open-source robot policies, including the autoregressive policy OpenVLA and the diffusion-based policies Octo and $\pi_{0}$. The obtained results in both simulation and real-world environments prove PCD’s flexibility and effectiveness, e.g., PCD enhances the state-of-the-art policy $\pi_{0}$ by 8% in the simulation environment and by 108% in the real-world environment.
@@ 1 Introduction
Figure 1: Robot policies tend to spuriously correlate task-irrelevant features with actions, compromising their ability to generalize to unseen scenarios. As observed, changing the light position from (a) to (b) and the drawer handle position from (a) to (c) results in 36% and 32% drops in the performance of the baseline policy OpenVLA [20], respectively. (d) Attention map. More results are in Section 4.4 and Appendix A.1.
@@
Contributions. Our contributions in this work are threefold. 1) We propose PCD, a simple, training-free, and easy-to-implement scheme to tackle the spurious correlation issue in robot policies. 2) PCD can be used as a plugin to enhance both autoregressive and diffusion-based policies, without requiring fine-tuning or access to pre-trained model weights. 3) Extensive experiments demonstrate PCD’s effectiveness and flexibility, consistently improving performance over three state-of-the-art policies across 15 manipulation tasks on multiple robotic platforms.
@@ 3.2 Policy Contrastive Decoding (PCD)
where $C$ is a normalization constant and a larger value of $\alpha\geq 0$ indicates a stronger amplification of differences between the two distributions. Particularly, $\alpha=0$ recovers the baseline policy. In such a manner, the obtained $\pi_{\theta}^{*}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})$ amplifies model predictions on object-relevant features in $\boldsymbol{u}$, thus exhibiting insensitivity to spurious features in $\boldsymbol{v}$, as illustrated in Fig. 2.
@@ 4 Experiments
In this section, we conduct extensive experiments in both simulation and real-world environments, on top of three state-of-the-art open-source robot policies to answer the following questions: {mdframed} [skipabove=4pt, innertopmargin=4pt, innerbottommargin=4pt] 1) Can PCD improve robot policies in both simulation and real-world environments?
@@ 4.4 Robustness to Various Kinds of Spurious Correlations
The central concept of our proposed PCD approach in this work is to overcome the adverse effects of spurious features on the decision-making of robot policies. So, what types of spurious correlations can PCD effectively address? To answer the question, we evaluate the robustness of the OpenVLA [20] and $\pi_{0}$ [3] policies to task scene variations in both simulation and real-world environments. As illustrated in Fig. 5, we report the performance of the two baseline policies integrated w/ or w/o our proposed PCD approach in different testing scenarios with unseen spatial relationships, brightness, distractors and table textures within visual observations. The results shown in the figure lead to the following observations. 1) Consistent with the previous experimental results, our PCD approach significantly improves the performance of the baseline policies on both simulation and real-world tasks (Orange line vs. Blue line). 2) Both baseline policies experience substantial drops in performance when testing scenarios are modified. For example, altering the brightness of visual observations results in 48% and 75% decreases in success rates for the two policies, respectively. 3) PCD consistently alleviates the performance degradation of the baseline policies across all unseen testing scenarios. Remarkably, after integrating PCD, the two baseline policies even performs better in 4/10 of unseen scenarios than in the original training scenarios. The results in the figure reveal that our proposed PCD approach can serve as a plugin to mitigate the adverse effects of various types of spurious correlations on the generalization of robot policies.
[…]